Find the area of region bounded between the parabola and line given below:

a. y2 = 2x and y = 2x.

b. y2 = 4x and y = 2x.

c. y2 = 4ax and y = x.

d. y2 = 2kx and y = mx.


1
Expert's answer
2021-04-29T14:25:09-0400

a) the points of intersection of the parabola "y^2= 2x" and the line y=2x are points for whivh;

"\sqrt{2x}= 2x \implies 2x=4x^2; 4x^2-2x=0"

"2x(2x-1)=0 \implies x=0, x=\frac 12 ...(i)"

now, "\int ^{\frac12}_0 (\sqrt{2x}-2x)dx= [\frac 23 \sqrt 2 x^{\frac32}-x^2]^{\frac12}_0"

"=[\frac 23 \sqrt2 (\frac 12)^{\frac 32}-(\frac 12)^2]=0"

"= \frac 1{12}" square units of length


b) the points of intersection of the curves are;"(2x)^2=4x \implies 4x^2 = 4x = 4x^2-4x=0; 4x(x-1)=0"

"x=0, x=1 ...(i)"

thus,

"area= \int^1_0 (2\sqrt x -2x)dx= [\frac 43 x^{\frac 32}- x^2]_0^1"

"= [\frac 43 (1)^{\frac 32}-(1)^2]-0= \frac 13" square units of length

c) the points of intersection of the parabola"y^2= 4ax" and the line y=x happen when

"2\sqrt {2x}=x \implies 4ax=x^2"

"\therefore x^2-4ax=0; x(x-4a)=0"

"x=0, x=4a ...(i) \forall a>0"

thus

"area= \int^{4a}_0 (2\sqrt{2x}-x)dx= [\frac 43 \sqrt a x^{\frac32} -\frac {x^2}2]^{4a}_0"

"=[\frac 43 \sqrt a (4a^{\frac 32} -\frac {(4a)^2}{2}]= \frac 83 a^2" square units of length

d) let k>0 and we may assume m>o without loss of generality (since for m<0, the region will be symmetric to that of m>0 but only below x-axis);

now, the points of intersection are where "(mx)^2 = 2kx \implies m^2 x^2 - 2kx=0; x(xm^2-2k)=0"

"x=0, x=\frac {2k}{m^2} ...(i)"

thus

"area = \int^{\frac {2k}{m^2}}_0 (\sqrt {2kx}-mx)dx= [\frac 23 \sqrt {2k}x^{\frac 32} -\frac {mx^2}{2} ]^{\frac {2k}{m^2}}_0"

"=[\frac 23 \sqrt{2k} (\frac {2k}{ m^2})^{\frac 32}- \frac m2(\frac {2k}{m^2})^2]=\frac {2k^2}{3m^2}" square units of length